OpenStudy (anonymous):

Heres the equation: h= 5.2sin2(pi)(t-4/12.4) +6.4 a. At what time, t, did the height reach 7m? Solve for t. 7=5.2sin2(pi)(t-4/12.4)+6.4 Please explain your answer.

OpenStudy (anonymous):

\[t = \sin^{-1}[ (0,6*6,2/5,2*\pi) +4\pi]\]